OpenStudy (anonymous):

which equation represents the line through the points (4,2) and is perpendicular to the equation represented by the equation y=-1/2x+3?

OpenStudy (abbles):

-1/2 is the slope of the current equation and you need to find the opposite reciprocal to find the equation of the line perpendicular to that.

OpenStudy (bakonloverk):

y-y1=m(x-x) y-2=-1/2(x-4) y-2=2x-2 +2 +2 y=2x